Three identical dies are rolled. The probability that the same number will appear on each of them is

Grade:12

1 Answers

bharat bajaj IIT Delhi
9 years ago
Total number of cases = 6 × 6 × 6 = 216
Same number can appear as (1, 1, 1), (2, 2, 2), (3, 3, 3) ... (6, 6, 6).
Hence no. of favourable cases = 6
Required probability = 6 / 216 = 1 / 36
Thanks
Bharat Bajaj
IIT Delhi
askiitians faculty